### v2

@@ Abstract
An approach to improve neural network interpretability is via clusterability, i.e., splitting a model into disjoint clusters that can be studied independently. We define a measure for clusterability and show that pre-trained models form highly enmeshed clusters via spectral graph clustering. We thus train models to be more modular using a “clusterability loss” function that encourages the formation of non-interacting clusters. Using automated interpretability techniques, we show that our method can help train models that are more modular and learn different, disjoint, and smaller circuits. We investigate CNNs trained on MNIST and CIFAR, small transformers trained on modular addition, and language models. Our approach provides a promising direction for training neural networks that learn simpler functions and are easier to interpret.
@@ 1 Introduction
Interpretability is an active area of research that aims to solve both high-stake deployment constraints for fairness and robustness (McClure et al., 2020) and AI safety and trustworthiness concerns (Bereska & Gavves, 2024). Several breakthroughs in the subdomain of mechanistic interpretability, have helped us understand the inner workings of deep networks, both via circuits (Wang et al., 2022, Olah et al., 2020, Elhage et al., 2021) and representation spaces (Zou et al., 2023, Bricken et al., 2023).
One approach to mechanistic interpretability is to identify features at the level of neurons, alternatively we can identify ‘subskills’ at the level of subnetworks or circuits. While sparse auto-encoders (SAEs) (Huben et al., 2024) extract a model’s features over different neurons, we nudge the model to divide its computation into disentangled components.
In this work, we split models into separate modules and interpret them in isolation. However, this is feasible only if the interaction, between these modules is minimal, i.e. if the clusterability is high. We introduce a metric to measure the amount of clusterability in neural network components, and attempt to make models modular and more interpretable during training by optimizing for this metric.
@@
Making models modular has two interpretability benefits: (a) modules may specialize in different sub-skills and behaviors, which would allow us to investigate and control them independently; and (b) computationally, interpretability methods such as patching Wang et al. (2022), pruning, and training sparse auto-encoders Cunningham et al. (2023) become more tractable.
We empirically show that when we train a model to be modular, the modules specialize in different ways. For models trained on CIFAR, we find in Figure 6 that modules specialize on different labels. For language models we find in Figure 7 that in models trained to be modular, performance often only requires a single module.
Applying mechanistic interpretability methods to large models (Kaplan et al., 2020) and complex behaviors has been a major hurdle for interpretability Lieberum et al. (2023), Golechha & Dao (2024) due to complicated computational subgraphs (circuits) and superposition Henighan et al. (2023), i.e., networks representing more features than they have neurons. We theoretically show in Section 7.2 that the latent space of a modular matrix has a much smaller feature-space, and empirically show for CIFAR in Section 6.2 that splitting models into separate components makes circuits smaller, making pruning and patching (Conmy et al., 2023) more tractable.
@@
  • We use automated interpretability methods to show that the clusters thus obtained make the model more interpretable by producing specialized clusters (Section 6.1).
@@
We also modify and test prior clusterability methods to split a trained neural network layer into clusters and show that the clusters thus found are highly enmeshed (not modular) and do not help our interpretability goals (Section 4), discuss how our training methodology, as described in Section 5.1, can help with building or fine-tuning more modular models, and discuss some mathematical implications in Section 7.
@@ 2 Related Work
metrics inspired by research in neuroscience incorporate transfer entropy (Novelli et al., 2019, Ursino et al., 2020) or spatial metrics (Liu et al., 2023a; b). Models are sometimes trained or pruned using modularity metrics (Patil et al., 2023). Tsitsulin et al. (2023) find that Graph Neural Networks node pooling methods do not work well to cluster graphs and introduce an unsupervised clustering method. Their method first uses a graph convolutional network to obtain soft clusters for each node and then optimizes this assignment based on the first modularity metric in Section 3. Salha-Galvan et al. (2022) add a regularization term for modularity in Graph Autoencoders to show performance gains. In this work, we focus on how modularity can help us train models to be more interpretable.
@@
is a form of modularity, where each module is an entire network. In our set-up each module is a component of a layer, i.e. the partitioning is at a lower level. Mixture of experts sometimes use routing policies to decide which expert to activate (e.g. based on features of the input). Gradient routing (Cloud et al., 2024) is a method that trains task-specialized experts by hard-coding what tasks each expert specializes on, one of their stated benefits is that this allows a user to selectively remove ability on a particular task. MONET (Park et al., 2024) trains monosemantic experts, which learn to route different types of inputs to specific experts during training, without human-specified task assignments. Instead, we encourage the emergence of task-specialized layer components during training.
@@
of neural network weights is typically done either using graph properties of the weights (structural) (Watanabe et al., 2018, Filan et al., 2021, Patil et al., 2023) or using correlations between neuron activations (functional) (Hod et al., 2022, Lange et al., 2022). MoEfication groups feedforward neurons in a pre-trained language model into clusters of ‘experts’ and at inference only activates the most clusters neurons (Zhang et al., 2022). In this paper, we consider both weight-based and gradient-based clustering, and show that they do not help with modularity across clusters.
@@
is an active field of research (Wang et al., 2022, Olah et al., 2020, Elhage et al., 2021, Conmy et al., 2023), which aims to uncover subnetworks that perform a specific functionality. Wortsman et al. (2020) use a randomly initialized, fixed base network and for each task find a subnetwork that achieves good performance on this task. These overlapping subnetworks can be thought of as circuits. In this work, we evaluate and optimize for structural connectedness or global functionality without considering specific functionalities. We evaluate the resulting modules by assessing the extent to which they have specialized on a class level, i.e. the extent to which they correspond to class-specific performance.
@@ 3 Introducing a Measure of Modularity
Non-differentiable metrics to measure modularity that rely on conditional entropy (Ursino et al., 2020), sampling, or discrete computation Veniat et al. (2021) have been introduced. While they capture the essence of modularity, we cannot use gradient descent to optimize for them during training. For unweighted graphs, ‘community structure’ is a commonly used modularity metric (Newman, 2006, Salha-Galvan et al., 2022, Tsitsulin et al., 2023, Bhowmick et al., 2024), which is based on the difference between an edge value (0 or 1) and the expected number of edges (see Appendix H.1 for an explicit formula).
@@
  • Mathematically, as discussed in Section 7, optimizing our metric leads to models learning simpler function spaces (polytopes) and fewer features (orthogonal directions) to perform a given task.
  • As we show in Section 5.2, optimizing for clusterability leads to different, disjoint, and smaller circuits (in the case of CIFAR-10) and other interpretability benefits.
@@ 4 Neural Networks trained on Cross-Entropy are not Modular
Our clustering method, called Bipartite Spectral Graph Clustering (BSGC) is shown in Algorithm 1. The similarity matrix for BSGC can be created by either the weights of the model or the accumulated gradients.
Weight-based BSGC.
Gradient-based BSGC.
In Figure 2, we see that as the number of clusters increases, the amount of clusterability decreases, and even at $k=2$ clusters, we get $C=0.6$ (random is 0.5, higher is better), which is substantial interference between clusters and hinders our interpretability goals. We show similar results for Pythia-70m Biderman et al. (2023) in Appendix G.
5 Methodology
Across a range of models and tasks, we optimize for modularity by training to maximize the clusterability (as defined in Section 3), and use various automated interpretability methods to evaluate the interpretability gains of the clusters thus obtained. This section details our methodology, and in Section 6 we share our results.
5.1 Training for Modularity
  1. 1.
    Train the original model to minimize cross-entropy for the first $t$ steps. This can help to form simpler features before we begin clustering and allow “winning tickets” (as defined in the lottery ticket hypothesis (Frankle & Carbin, 2018)) to emerge. In practice, we found that our results do not vary with $t$, and default to $t=0$ for most results.
  2. 2.
  3. 3.
  4. 4.
We experiment with vanilla MLPs on MNIST (Deng, 2012), and CNNs on the CIFAR-10 dataset (Krizhevsky et al., 2009), transformers on modular addition, and fine-tuning of large language models (LLMs). We compare the performance, clusterability, and interpretability gains of our clustered models against models trained without the clusterability loss. Our results are available in Section 6 and all the hyperparameters we use are given in Appendix A.
Figure 3: Clusterability against the train and test losses of a 1-layer transformer model trained modularly on modular addition. For this task, we show that a model can be trained to completely solve the task while making every model component completely clusterable.
For modular addition, we follow the training pipeline of Nanda et al. (2023) to train a 1-layer transformer model and increase the training data to reduce “grokking”, or the delayed generalization effect they studied. Instead, we focus on training the model with the clusterability loss, and show in Figure 3 that we can train the model to completely solve the task while making every model component completely clusterable. See Appendix D for an exploration of the interpretability gains from modularity in this setting.
For language models, we fine-tune our models on a subset of the Wikipedia data Merity et al. (2016), while making the MLP input and output matrices to be modular. For individual layers, we define the “maximum clusterability” of a layer to be the maximum amount of clusterability that can be achieved without loss of performance. We measure this by training the model with the clusterability loss until the clusterability stops improving or the performance starts degrading. We show the maximum clusterability for the MLP input and output matrices of each layer in Figure 4.
Figure 4: Maximum clusterability achievable for each layer’s MLP input and output matrices (independently) for GPT2-small without degradation in performance. We found the MLP input matrices to have higher maximum clusterability compared to the MLP output matrices, except in the first layer.
Figure 5: Combined clusterability of the MLP input of all layers in GPT2-small and the cross-entropy training loss on a subset of the Wikipedia dataset during training for modularity.
We show the training plots for modular training of all MLP_in weight matrices with the cross-entropy training loss in Figure 5. We checkpoint the model after the first epoch to avoid overfitting and study the model for interpretability gains, comparing it with a model fine-tuned on the same data for the same number of epochs but without the clusterability loss.
5.2 Evaluating Interpretability Gains
We use various automated interpretability metrics to evaluate our clustered models and their interpretability gains. In this section, we define these metrics: class-wise accuracy (with and without each individual cluster), and Effective Circuit Size (ECS) for each label. (See Section 6 for our results on these metrics.)
5.2.1 Interventions
  1. 1.
  2. 2.
5.2.2 Effective Circuit Size (ECS)
We use automated pruning to recursively remove edges that do not contribute (see Automated Circuit DisCovery (Conmy et al., 2023)) to extract the ‘effective circuit’ for each label and define the ratio of the number of parameters in it to the number of parameters in the whole model to be the effective circuit size (ECS) for that label and model. A lower effective circuit size indicates that the model has learned a smaller task-specific circuit, and reduction in ECS is a proxy for interpretability gains. Since recursively removing edges is expensive, we only show results on simpler models, and leave more efficient pruning techniques for future work.
6 Results
6.1.1 Clusters Specialize in Class-level Features for MLPs and CNNs
Figure 6: Class-wise accuracy for each label with clusters turned ON (top) and OFF (bottom). Note that individual clusters learn near-complete circuits for various labels, such as Cluster 00 for SHIP and Cluster $1$ for TRUCK. The figure on the top shows that a cluster is “sufficient” to predict a given class, while a dip in the accuracy in figure on the bottom shows that it is “necessary”. For instance, note that Cluster $3$ is both necessary and sufficient for completely identifying a “deer”.
Figure 6 compares class-wise accuracy of the model on the CIFAR-10 dataset with individual clusters turned OFF and ON respectively in a clustered model with an accuracy $>95\%$ for each label. We find individual clusters learning near-complete circuits for various labels for both kinds of interventions described in Section 5.2.1. In other words, clusters specialize on certain labels.
Figure 7: Fraction of samples (left) that need $k$ clusters for correct prediction, for $k\in[1,2,3,4]$ for a clustered model with $4$ clusters. Note that in a modular model, most datapoints can be solved by $1-2$ clusters, while $3-4$ are required for a large fraction in a non-modular model fine-tuned on the same data. Figure (b) on the right shows the number of datapoints each cluster contributes to, indicating that a modular model reduces the number of clusters involved and keeps them all similarly useful.
6.1.2 Task Specialization in Modular Language Model (LLM) Fine-tuning
We also do an ON-intervention investigation on GPT2-small fine-tuned on wiki data. In Figure 7(a), we see the fraction of samples (left) that need $k$ clusters for correct prediction, for $k\in[1,2,3,4]$ for a clustered model with $4$ clusters. Note that in a modular model, most datapoints can be solved by $1-2$ clusters, while $3-4$ are required for a large fraction in a non-modular model fine-tuned on the same data. See Appendix F for a more elaborate explanation of these values.
Figure 7(b) on the right shows the number of datapoints each cluster contributes to, indicating that a modular model reduces the number of clusters involved and keeps them all similarly useful.
6.2 Modularity Encourages Smaller Circuits
Figure 8 compares the Effective Circuit Size (ECS) for each label for the clustered and unclustered models on CIFAR-10 (see Section 5). We show that an unclustered model has, on average, $61.25\%$ more parameters in its effective circuits. In Appendix B, we share similar results for the MNIST dataset (Deng, 2012).
7 Theoretical Analysis of the Effects of Modularity
7.1 Polytopes in a Modular Network
We show that the partitioning of neural networks into polytopes on which the network is linear (Sudjianto et al., 2020) becomes coarser when we make a model modular. In other words, the simplicity of the function space of a network increases by adding modularity constraints. In this section, we restrict to fully connected ReLU feed-forward neural networks.
Definition 7.1 (ReLU Network).
$\mathcal{N}(x)=W^{(L+1)}\chi^{(L)}+b^{(L+1)}.$
The number of neurons in each layer is given by a vector $\mathbf{N}=[n_{1},n_{2},\ldots,n_{L}]$, and all activations are in the positive reals, i.e. $\chi^{(l)}\in\mathbb{R}^{n_{l}}_{\geq 0}$ for all $l\in\{1,\ldots,L\}=[L]$. We stress the dependence on $x$ by writing $\chi^{(l)}(x)$.
ReLU networks can be described as linear models that are applied to certain regions of the input space (Sudjianto et al., 2020). These regions partition the input space $\mathbb{R}^{n}$.
Proposition 7.2.
$f|_{\omega}(x)=\alpha_{\omega}^{T}x+\beta_{\omega}.$
Moreover, each part is a polytope, given by the intersection of a collection of half-spaces. We write the minimum set of half-space conditions that can be used to specify the entire partition as $H_{1},\ldots,H_{k}$, where each $H_{i}$ is given by the set of all $x\in\mathbb{R}^{n}$ such that, for for $h_{i,j}\in\mathbb{R},i\in[k],j\in[n]$:
$h_{i,1}x_{1}+h_{i,2}x_{2}+\ldots+h_{i,n}x_{n}>c_{i}.$
In some sense, the number of polytopes gives a measure of the granularity or refinement of a network. Given a network with neurons $\mathbf{N}=[n_{1},n_{2},\ldots,n_{L}]$, there are at most $2^{n_{1}}\times 2^{n_{2}}\times\ldots\times 2^{n_{L}}$ polytopes. Now suppose that the weight matrix $W^{(l)}$ in layer $\chi^{(l)}$ is modular and contains $k$ modules, meaning that we can divide $n_{l-1}$ and $n_{l}$ into $k$ sets $[n_{l-1}^{1},\ldots,n_{l-1}^{k}]$ and $[n_{l}^{1},\ldots,n_{l}^{k}]$ such that the only non-zero weights in $W^{(l)}$ are between $n_{l-1}^{i}$ and $n_{l}^{i}$. In this case the number of half-space conditions generated by $\chi^{(l-1)}$ and $\chi^{(l)}$ is no longer $2^{n_{l-1}}\times 2^{n_{l}}$, but it now is $2^{n_{l-1}}\times 2^{n_{l}^{1}}+\ldots+2^{n_{l-1}}\times 2^{n_{l}^{k}}$, which is substantially fewer. This means that modularity increases the simplicity of the function-space of a network at the cost of expressivity.
7.2 Concept Representations in Activation Space
Proposition 7.3 (Johnson-Lindenstrauss lemma (Johnson & Lindenstrauss, 1984)).
Let $10<\epsilon<$ 1 and let $X$ be a set of $m$ points in $\mathbb{R}^{N}$ and $n>\frac{8(ln(m))}{\epsilon^{2}}$, then there exists a linear map $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}^{n}$ such that for all $u,v\in X$ we have
Then the number of nearly orthogonal points that can be represented in layer $n_{l}$ becomes ${}e^{n_{l}^{1}}+\ldots+e^{n_{l}^{k}}$ as opposed to ${}e^{{n_{l}^{1}}+\ldots+{n_{l}^{k}}}$. This means that as a result of modularizing the weight matrix $W^{(l)}$, the model learns to perform computation by using features from an exponentially smaller set of subspaces. Sparse Autoencoders (SAEs) Cunningham et al. (2023) extract linear subspaces as latents for human-interpretable features, and modularity helps reduce the search space for SAEs exponentially. Training SAEs on modular models and clusters is an interesting direction for future work.
8 Conclusion
We show that a simple regularizer is effective at splitting a neural network layer into simpler, more interpretable clusters. We show that the average circuit size and the search space improve with more clusters without a decrease in overall performance (for the models and tasks that we investigated).
Future work could study the Pareto frontier (clusterability versus performance) as we scale to harder tasks and larger models. We are also interested in using our insights to train and align language models with modularity and see if it leads to better control for harmful behaviors. We hope that modularity can help with a number of mechanistic interpretability goals, and a scaled-up exploration into this (such as clustering attention heads and training SAEs) is an interesting direction for future work.
9 Acknowledgments and Disclosure of Funding
We would especially like to thank Dylan Cope, Daniel Filan, Sandy Tanwisuth, Siddhesh Pawar, Niels uit de Bos, Matthew Wearden, and Henning Bartsch for valuable discussions, feedback, and support. SG,NS would like to thank the ML Alignment Theory & Scholars (MATS) Program, the organizers, funders, and staff. SG was supported by independent research grants from AI Safety Support and the Long-Term Future Fund. This research was supported by the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Hyperparameters

### v3

@@ Abstract
An approach to improve neural network interpretability is via clusterability, i.e., splitting a model into disjoint clusters that can be studied independently. We define a measure for clusterability and show that pre-trained models form highly enmeshed clusters via spectral graph clustering. We thus train models to be more modular using a “clusterability loss” function that encourages the formation of non-interacting clusters. We then investigate the emerging properties of these highly clustered models. We find our trained clustered models do not exhibit more task specialization, but do form smaller circuits. We investigate CNNs trained on MNIST and CIFAR, small transformers trained on modular addition, and GPT-2 and Pythia on the Wiki dataset, and Gemma on a Chemistry dataset. This investigation shows what to expect from clustered models.
@@ 1 Introduction
Interpretability is an active area of research that aims to solve both high-stake deployment constraints for fairness and robustness (McClure et al. 2020) and AI safety and trustworthiness concerns (Bereska and Gavves 2024). Several breakthroughs in the subdomain of mechanistic interpretability, have helped us understand the inner workings of deep networks, both via circuits (Wang et al. 2022; Olah et al. 2020; Elhage et al. 2021) and representation spaces (Zou et al. 2023; Bricken et al. 2023).
One approach to mechanistic interpretability is to identify features at the level of neurons, alternatively we can identify ‘subskills’ at the level of subnetworks or circuits. While sparse auto-encoders (SAEs) (Huben et al. 2024) extract a model’s features over different neurons, we attempt to split a model’s computation into separate modules and interpret them in isolation. However, this is feasible only if the interaction, between these modules is minimal, i.e. if the clusterability is high. We introduce a metric to measure the amount of clusterability in neural network components, and attempt to make models modular and more interpretable during training by optimizing for this metric.
@@
We empirically show that when we train a model to be modular, the modules do not specialize any more than non-modular models on the tasks that we test. For models trained on CIFAR, we find in Figure 6(a) that modules specialize on different labels, but so do non-modular models in Figure 6(b). For Gemma trained on a chemistry dataset, we find hardly any module specialization for either modular or non-modular models, see Figure 19. For GPT-2 and Pythia we find in Figure 8 that both modular and non-modular modules, are typically unable to solve a datapoint in isolation. In other words, we do not find any task specialization of modules. There may still be task specialization, on a different level of granularity than the ones we investigated.
Applying mechanistic interpretability methods to large models (Kaplan et al. 2020) and complex behaviors has been a major hurdle for interpretability (Lieberum et al. 2023; Golechha and Dao 2024) due to complicated computational subgraphs (circuits) and superposition (Henighan et al. 2023), i.e., networks representing more features than they have neurons. We theoretically show in Section 7.2 that the latent space of a modular matrix has a much smaller feature-space, and empirically show for CIFAR in Section 6.3 that splitting models into separate components makes circuits smaller, making pruning and patching (Conmy et al. 2023) more tractable.
@@
  • We use automated interpretability methods to show that the clusters thus obtained do not obviously produce specialized clusters (Section 6.1).
@@
We also modify and test prior clusterability methods (Filan et al. 2021) to split a trained neural network layer into clusters and show that the clusters thus found are highly enmeshed (not modular) in Section 4, discuss how our training methodology, as described in Section 5, can help with building or fine-tuning more modular models, and discuss some mathematical implications in Section 7.
@@ 2 Related Work
metrics inspired by research in neuroscience incorporate transfer entropy (Novelli et al. 2019; Ursino, Ricci, and Magosso 2020) or spatial metrics (Liu, Gan, and Tegmark 2023; Liu et al. 2023). Liu et al. (2023) incorporate neuron swapping, which is a discrete search over optimal neuron ordering. This swapping requires $\mathcal{O}(nkL)$ computations, where $n$ is the number of neurons going into a layer, $k<n$ is the number of neurons being sorted, and $L$ is the number of layers. Since swapping is expensive it is only performed every so many training steps. Models are sometimes trained or pruned using modularity metrics (Patil, Michael, and Dovrolis 2023). Tsitsulin et al. (2023) find that Graph Neural Networks node pooling methods do not work well to cluster graphs and introduce an unsupervised clustering method. Their method first uses a graph convolutional network to obtain soft clusters for each node and then optimizes this assignment based on the first modularity metric in Section 3. Salha-Galvan et al. (2022) add a regularization term for modularity in Graph Autoencoders to show performance gains. In this work, we investigate the relationship between modularity and interpretability.
@@
is a form of modularity, where each module is an entire network. In our set-up each module is a component of a layer, i.e. the partitioning is at a lower level. Mixture of experts sometimes use routing policies to decide which expert to activate (e.g. based on features of the input). Gradient routing (Cloud et al. 2024) is a method that trains task-specialized experts by hard-coding what tasks each expert specializes on, one of their stated benefits is that this allows a user to selectively remove ability on a particular task. MONET (Park et al. 2024) trains monosemantic experts, which learn to route different types of inputs to specific experts during training, without human-specified task assignments. We investigate the emergence of task-specialized layer components during training.
@@
of neural network weights is typically done either using graph properties of the weights (structural) (Watanabe, Hiramatsu, and Kashino 2018; Filan et al. 2021; Patil, Michael, and Dovrolis 2023) or using correlations between neuron activations (functional) (Hod et al. 2022; Lange, Rolnick, and Kording 2022). MoEfication groups feedforward neurons in a pre-trained language model into clusters of ‘experts’ and at inference only activates the most clustered neurons (Zhang et al. 2022). In this paper, we consider both weight-based and gradient-based clustering, and show that they do not help with modularity across clusters.
@@
is an active field of research (Wang et al. 2022; Olah et al. 2020; Elhage et al. 2021; Conmy et al. 2023), which aims to uncover subnetworks that perform a specific functionality. Wortsman et al. (2020) use a randomly initialized, fixed base network and for each task find a subnetwork that achieves good performance on this task. These overlapping subnetworks can be thought of as circuits. In this work, we investigate if modularity can help with reducing circuit size.
@@ 3 Introducing a Measure of Modularity
Non-differentiable metrics to measure modularity that rely on conditional entropy (Ursino, Ricci, and Magosso 2020), sampling, or discrete computation (Veniat, Denoyer, and Ranzato 2021) have been introduced. While they capture the essence of modularity, we cannot use gradient descent to optimize for them during training. For unweighted graphs, ‘community structure’ is a commonly used modularity metric (Newman 2006; Salha-Galvan et al. 2022; Tsitsulin et al. 2023; Bhowmick et al. 2024), which is based on the difference between an edge value (0 or 1) and the expected number of edges (see Appendix J.1 for an explicit formula).
@@
  • In this paper we apply the modularity metric to a single weight matrix, but we can easily generalize the approach to multiple weight matrices. To create modules that are multiple layers deep, one should ensure the neurons in a single cluster in the output layer of the first matrix match the neurons in the input layer of the second matrix and so on.
  • Mathematically, as discussed in Section 7, optimizing for our metric leads to models learning simpler function spaces (fewer polytopes) and fewer features (orthogonal directions) to perform a given task, which can be computationally beneficial when model interpretability or explainability is important.
  • As we show in Section 6.3, optimizing for clusterability leads to smaller circuits (in the case of CIFAR-10).
@@ 4 Neural Networks trained on Cross-Entropy are not Modular
Our clustering method, called Bipartite Spectral Graph Clustering (BSGC) is shown in Appendix A Algorithm 1. The similarity matrix for BSGC can be created by either the weights of the model or the accumulated gradients.
Weight-based BSGC.
Gradient-based BSGC.
In Figure 2, we see that as the number of clusters increases, the amount of clusterability decreases, and even at $k=2$ clusters, we get $C=0.6$ (random is 0.5, higher is better), which is substantial interference between clusters. We show similar results for Pythia-70m (Biderman et al. 2023a) in Appendix H.
In the remainder of this paper, we arbitrarily choose $k=4$. In general, we would recommend choosing $k$ based on task specific considerations, possibly doing a sweep.
5 Training for Modularity
  1. 1.
    Train the original model to minimize cross-entropy for the first $t$ steps. This can help to form simpler features before we begin clustering and allow “winning tickets” (as defined in the lottery ticket hypothesis (Frankle and Carbin 2018)) to emerge. In practice, we found that our results do not vary with $t$, and default to $t=0$ for most results.
  2. 2.
  3. 3.
  4. 4.
Figure 3: Clusterability vs. train and test losses of a 1-layer transformer trained modularly on modular addition.
Figure 4: Maximum clusterability of MLP input/output matrices in GPT2-small without degrading performance.
We experiment with vanilla MLPs on MNIST (Deng 2012), and CNNs on the CIFAR-10 dataset (Krizhevsky, Hinton et al. 2009), transformers on modular addition, and fine-tuning of large language models (LLMs). We compare the performance, clusterability, and interpretability gains of our clustered models against models trained without the clusterability loss. Our results are available in Section 6 and all the hyperparameters we use are given in Appendix B.
For modular addition, we follow the training pipeline of Nanda et al. (2023) to train a 1-layer transformer model and increase the training data to reduce “grokking”, or the delayed generalization effect they studied. Instead, we focus on training the model with the clusterability loss, and show in Figure 4 that we can train the model to completely solve the task while making every model component completely clusterable. See Appendix E for an exploration of the interpretability gains from modularity in this setting.
For language models, we fine-tune GPT-2 (Lee and Hsiang 2019) and Pythia (Biderman et al. 2023b) on a subset of the Wikipedia data (Merity et al. 2016), while making the MLP input and output matrices to be modular. We also fine-tune Gemma (Team et al. 2024) on a Chemistry dataset (XythicK 2025). For individual layers, we define the “maximum clusterability” of a layer to be the maximum amount of clusterability that can be achieved without loss of performance. We measure this by training the model with the clusterability loss until the clusterability stops improving or the performance starts degrading. We show the maximum clusterability for the MLP input and output matrices of each layer in Figure 4.
We show the training plots for modular training of all MLP_in weight matrices with the cross-entropy training loss in Figure 13. We checkpoint the model after the first epoch to avoid overfitting and study the model for interpretability gains, comparing it with a model fine-tuned on the same data for the same number of epochs but without the clusterability loss. The jumps in the cross-entropy loss are due to the model learning from a small enough batch to affect performance on the other datapoints.
6 Properties of Modular Models
In this section we do two investigations into the extent to which the modules of a modular model specialize for specific tasks. We test task-specialization for CNNs on CIFAR in Section 6.1 and find that both modular and non-modular models specialize into specific classes. In contrast, in Section 6.1 we also test whether modules specialize to specific categories of chemistry problems, and find that neither modular nor non-modular models do. We hypothesize that whether modules specialize or not is related to whether features are localized or not.
Modules Specialize in Class-level Features for Both Modular and Non-Modular Models
Figure 5: Class-wise accuracy for each label with clusters turned ON (top) and OFF (bottom). Note that individual clusters learn near-complete circuits for various labels, such as Cluster 0 in the clustered model for SHIP and Cluster 111 for TRUCK for the modular model. The figure on the top shows that a cluster is “sufficient” to predict a given class, while a dip in the accuracy in figure on the bottom shows that it is “necessary”.
Figure 6: Intervention results on a Chemistry dataset with a clustered Gemma-1B model with 444 clusters denoted by the four colors. Note that the non-modular model performs better than the modular model.
[…]
Modules Do Not Specialize in Class-level Features for Both Modular and Non-Modular LLMs
  1. 1.
  2. 2.
In Figure 19, we show that optimizing the Gemma model for modularity on a Chemistry question-answering task with semantic categories (XythicK 2025) does not yield cluster-level specialization. We also find the performance to be worse for the modular model.
GPT-2 Specializes on Specific Datapoints to Some Extent
We also do an ON-intervention investigation on GPT2-small fine-tuned on wiki data. In Figure 7, we see the fraction of samples for which there are exactly $k$ clusters that are individually able to correctly predict the answer. Note that in a modular model, most datapoints can only be solved by 121-21 - 2 clusters, while 343-43 - 4 modules can interchangably solve for a large fraction of datapoints in a non-modular model fine-tuned on the same data. See Appendix G for a more elaborate explanation of these values.
Figure 7: Fraction of samples for which there are exactly $k$ clusters that are individually able to correctly predict the answer. Here $k\in[1,2,3,4]$ for a clustered model with 444 clusters.
6.2 Individual Competency of Modules in LLMs
We also check whether modules are able to independently solve datapoints in Figure 8. Here we check for how many datapoints there were no modules that could solve the datapoint. We find that for GPT-2 modular models had much more difficulty with predicting datapoints using a single module, as the number of incorrectly predicted samples was 80%. For all other settings, modular and non-modular models behaved similarly, with modular modules being ever so slightly worse at predicting samples (using one or three modules).
(a) One module turned on in layer 5
(b) Three modules turned on in layer 5
Figure 8: Turning modules on or off in LLMs trained to be modular, or in LLMs where we arbitrarily select modules.
6.3 Modularity Encourages Smaller Circuits
We use automated pruning to recursively remove edges that do not contribute (see Automated Circuit DisCovery (Conmy et al. 2023)) to extract the ‘effective circuit’ for each label and define the ratio of the number of parameters in it to the number of parameters in the whole model to be the effective circuit size (ECS) for that label and model. A lower effective circuit size indicates that the model has learned a smaller task-specific circuit, and reduction in ECS is a proxy for interpretability gains. Since recursively removing edges is expensive, we only show results on simpler models, and leave more efficient pruning techniques for future work.
[…]
Our clusterability measure regularizes against connections between modules. In the 271424 parameter CNN model we use, there are 1024 such connections. So even if we all of these connections were zero, that would only be a reduction of 0.37% of the total number of parameters. This means that the reduction in parameters in effective circuits is unlikely due to regularization leading to a reduction in non-zero parameters, and instead due to underlying restructuring of the circuits.
7 Theoretical Analysis of the Effects of Modularity
7.1 Polytopes in a Modular Network
We show that the partitioning of neural networks into polytopes on which the network is linear (Sudjianto et al. 2020) becomes coarser when we make a model modular. In other words, the simplicity of the function space of a network increases by adding modularity constraints. In this section, we restrict to fully connected ReLU feed-forward networks.
Definition 7.1 (ReLU Network).
$\mathcal{N}(x)=W^{(L+1)}\chi^{(L)}+b^{(L+1)}.$
The number of neurons in each layer is given by a vector $\mathbf{N}=[n_{1},n_{2},\ldots,n_{L}]$, and all activations are in the positive reals, i.e. $\chi^{(l)}\in\mathbb{R}^{n_{l}}_{\geq 0}$ for all $l\in\{1,\ldots,L\}=[L]$.
ReLU networks can be described as a collection of linear models that are applied to certain regions of the input space (Sudjianto et al. 2020). These regions partition the input space $\mathbb{R}^{n}$.
Proposition 7.2.
$f|_{\omega}(x)=\alpha_{\omega}^{T}x+\beta_{\omega}.$
Moreover, each part is a polytope, given by the intersection of a collection of half-spaces.
In some sense, the number of polytopes gives a measure of the granularity or refinement of a network. Given a network with neurons $\mathbf{N}=[n_{1},n_{2},\ldots,n_{L}]$, there are at most $2^{n_{1}}\times 2^{n_{2}}\times\ldots\times 2^{n_{L}}$ polytopes. Now suppose that the weight matrix $W^{(l)}$ in layer $\chi^{(l)}$ is modular and contains $k$ modules, meaning that we can divide $n_{l-1}$ and $n_{l}$ into $k$ sets $[n_{l-1}^{1},\ldots,n_{l-1}^{k}]$ and $[n_{l}^{1},\ldots,n_{l}^{k}]$ such that the only non-zero weights in $W^{(l)}$ are between $n_{l-1}^{i}$ and $n_{l}^{i}$. In this case the number of half-space conditions generated by $\chi^{(l-1)}$ and $\chi^{(l)}$ is no longer $2^{n_{l-1}}\times 2^{n_{l}}$, but it now is $2^{n_{l-1}}\times 2^{n_{l}^{1}}+\ldots+2^{n_{l-1}}\times 2^{n_{l}^{k}}$, which is substantially fewer. This means that modularity increases the simplicity of the function-space of a network, which comes at the cost of expressivity.
7.2 Concept Representations in Activation Space
Proposition 7.3 (Johnson and Lindenstrauss (1984)).
Let $0<\epsilon<1$ and let $X$ be a set of $m$ points in $\mathbb{R}^{N}$ and $n>\frac{8(ln(m))}{\epsilon^{2}}$, then there exists a linear map $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}^{n}$ such that for all $u,v\in X$ we have
Then the number of nearly orthogonal points that can be represented in layer $n_{l}$ becomes $e^{n_{l}^{1}}+\ldots+e^{n_{l}^{k}}$ as opposed to $e^{{n_{l}^{1}}+\ldots+{n_{l}^{k}}}$. This means that as a result of modularizing the weight matrix $W^{(l)}$, the model learns to perform computation by using features from an exponentially smaller set of subspaces. Sparse Autoencoders (SAEs) (Cunningham et al. 2023) extract linear subspaces as latents for human-interpretable features, and modularity helps reduce the search space for SAEs exponentially. Training SAEs on modular models and modules is an interesting direction for future work.
8 Conclusion
We show that a simple regularizer can split a neural network’s layer into non-interacting modules, but these modules do not yield the interpretability benefits that were expected. We show that the average circuit size and the search space improve with more modules without a decrease in overall performance (for the models and tasks that we investigated).
Future work could investigate if there are alternative regularizers that do lead to clear interpretability benefits, and task specialization on harmful or desired tasks.
One possible hypothesis why our modularity metric does not work is that it operates on the wrong level of granularity, and further exploration into this would be an important future direction. To resolve this, future work could build an automated subtask detection mechanism, whereby the subtasks do not have to be specified beforehand.
Appendix A Algorithm
@@
Appendix B Hyperparameters